OpenStudy (anonymous):

What is the solution to the equation 1 over the square root of 8 = 4(m − 3)? m = −15/ 4 m = −11/4 m = 5/4 m = 9/4

OpenStudy (anonymous):

@ilikemath50

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

\(\dfrac{1}{\sqrt8}=4(m − 3) \) ?

OpenStudy (anonymous):

yes, but 4^(m-3)

OpenStudy (ikram002p):

\(\Huge \dfrac{1}{\sqrt8}=4^{(m − 3)}\) \(\Huge 2^{\frac{-3}{2}} =2^{2(m − 3)}\) \(\Huge {\frac{-3}{2}} = {2(m − 3)}\) solve for m

OpenStudy (ikram002p):

2m=-3/2+6 m=9/4

OpenStudy (anonymous):

thank you

OpenStudy (ikram002p):

np :)